OpenStudy (anonymous):

A bag has 4 green marbles, 3 red marbles, and 3 yellow marbles. What is the probability that you pick a red marble, do not replace it, and pick a green marble?

OpenStudy (amistre64):

how many marbles are there?

OpenStudy (anonymous):

10

OpenStudy (amistre64):

good, and how many are red?

OpenStudy (anonymous):

3

OpenStudy (amistre64):

great, so the prob of picking a red marble first is: 3/10 that leaves us 9 marbles left. how many that are left are green?

OpenStudy (anonymous):

4

OpenStudy (amistre64):

thats right, so the prob of picking a green marble next is 4/9 lets multiply the two results

OpenStudy (anonymous):

3/10 * 4/9 is. cross multiply?

OpenStudy (amistre64):

cross mulitply is a method for something else ... ratios when you multiply 2 fractions, just go straight across

OpenStudy (amistre64):

\[\frac ab \frac xy=\frac{ax}{by}\]

OpenStudy (anonymous):

2/15 then. thanks!

OpenStudy (amistre64):

your welcome